Let f(x)=|x-5| and g(x)=1/x
Write the given function as a composition function
using f and g.

H(x)=|1/x-(5)|
1
Expert's answer
2012-12-25T12:00:03-0500
H(x)=| 1/x -5 | let t=1/x then
H(x)=| t-5 |=f( t ) but t=1/x=g(x)
so we have H(x)=f(g(x))